# Optimal Best-Arm Identification under Fixed Confidence with Multiple Optima

### v1

@@ Abstract
We study the problem of best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, with a particular focus on instances that admit multiple optimal arms. While the Track-and-Stop algorithm of Garivier and Kaufmann (2016) is widely conjectured to be instance-optimal, its performance in the presence of multiple optima has remained insufficiently understood. In this work, we revisit the Track-and-Stop strategy and propose a modified stopping rule that ensures instance-optimality even when the set of optimal arms is not a singleton. Our analysis introduces a new information-theoretic lower bound that explicitly accounts for multiple optimal arms, and we demonstrate that our stopping rule tightly matches this bound.
Index Terms:
Multi-Armed Bandits, Best-Arm Identification, Sample Complexity.
@@ I Introduction
The multi-armed bandit (MAB) problem is a classic framework for sequential decision making under uncertainty, where a learner interacts with a set of arms representing unknown reward distributions. The objective is to gather information about these arms to make an informed decision.
In the best-arm identification (BAI) problem under the fixed-confidence setting, the learner’s goal is to identify an arm whose mean is maximal (or nearly so) with a prescribed confidence level, using as few samples as possible. This setting has broad applications in areas such as clinical trials, A/B testing, hyperparameter tuning, and recommendation systems.
While much of the prior work on BAI focuses on the case where there is a unique best arm, many practical scenarios involve multiple optimal arms—that is, more than one arm may achieve the highest expected reward. This work addresses a fundamental gap in the literature by developing a theoretical and algorithmic framework that achieves optimal sample complexity in the presence of multiple optimal arms.
@@ I-A Motivation
In best-arm identification problems, the primary goal is to identify an optimal arm from a set of stochastic options using as few samples as possible while maintaining a fixed confidence level. This problem is fundamental to sequential decision-making and has broad applicability in domains such as clinical trials, A/B testing, and hyperparameter tuning. While much of the existing work focuses on settings with a unique best arm, many real-world applications naturally involve multiple arms achieving the same maximal expected reward. This motivates the need to study best-arm identification in the more general scenario where the set of optimal arms is not a singleton.
Among algorithms proposed for fixed-confidence best-arm identification, the Track-and-Stop algorithm stands out due to its principled design based on likelihood ratio statistics and its strong empirical performance. Although it is widely believed to be near-optimal, especially in the case of a single best arm, a complete theoretical justification of its optimality—particularly in instances with multiple optimal arms—has remained an open question.
In this work, we address this gap by modifying the stopping rule of the Track-and-Stop algorithm and proving that the resulting algorithm achieves instance-optimal sample complexity in the fixed-confidence setting, even in the presence of multiple optimal arms. Our analysis introduces a refined information-theoretic lower bound tailored to this general setting and shows that the proposed stopping condition tightly matches this bound. This result provides the first formal guarantee of optimality for Track-and-Stop with multiple optima and offers new insights into the design of exploration strategies in stochastic bandit problems.
@@ I-B Related Work
The problem of best-arm identification (BAI) in stochastic multi-armed bandits has been a central focus in the bandit literature for over a decade, particularly under the fixed-confidence and fixed-budget settings. In the fixed-confidence framework, the goal is to identify an arm with the highest expected reward with probability at least $1-\delta$, while minimizing the expected number of samples (even2006action, ; garivier2016optimal, ). The fixed-budget setting, on the other hand, fixes the total sampling budget and aims to minimize the error probability (gabillon2012best, ).
@@
Early foundational work in fixed-confidence BAI introduced elimination strategies and stopping criteria that guarantee PAC (Probably Approximately Correct) bounds (even2006action, ). These were later complemented by information-theoretic approaches that established matching lower and upper bounds on sample complexity under the assumption of a unique best arm (garivier2016optimal, ; garivier2016track, ). Such analyses rely on likelihood ratio statistics and characterize the complexity of the problem through Kullback-Leibler (KL) divergences between the reward distributions of arms. Algorithms such as Track-and-Stop (garivier2016optimal, ) and lil’UCB (jamieson2014lil, ) leverage these theoretical insights to achieve near-optimal sample efficiency.
Algorithms like LUCB (kalyanakrishnan2012pac, ) and lil’UCB adapt confidence bounds to balance exploration across arms efficiently, achieving PAC guarantees with sample complexities that depend on the inverse squared gaps between arm means. Despite their success, these methods generally rely on the uniqueness of the best arm, and their performance can degrade when multiple arms share the highest mean reward.
In many practical applications—such as clinical trials, recommendation systems, and A/B testing—multiple arms may be equally optimal due to noise or inherent problem structure. The presence of multiple optimal arms poses unique challenges because standard algorithms continue to sample these tied arms excessively to differentiate among them, leading to unnecessary increases in sample complexity.
The literature addressing multiple optimal arms has mostly focused on related but distinct problems such as top-$k$ arm identification or best subset selection (gabillon2012best, ; chen2014combinatorial, ; jamieson2014bandit, ). These works aim to identify the entire set of best arms with high confidence, often requiring more samples than necessary if the goal is only to find any one best arm. Some recent studies have begun to recognize this inefficiency; for instance, (qin2022gap, ) propose gap-free algorithms that adapt automatically to the presence of zero gaps (i.e., ties) and reduce redundant exploration. Similarly, (combes2021minimal, ) analyze minimal exploration strategies and highlight how ignoring ties can cause suboptimal sampling strategies.
More recent advances by (hu2023almost, ) propose algorithms explicitly designed to exploit the existence of multiple optimal arms. These algorithms feature stopping rules and sampling strategies that avoid unnecessary comparisons between tied arms and achieve nearly instance-optimal sample complexity under fixed confidence. Their approach shows significant improvements over classical methods when multiple arms are truly optimal.
From a theoretical standpoint, classical lower bounds on sample complexity assume uniqueness of the best arm (garivier2016optimal, ). The multiple-optimal-arm scenario requires refined lower bounds that reflect the complexity of identifying any one of the optimal arms, demonstrating how sampling effort should be distributed among tied arms rather than wasted on differentiating them. This nuanced understanding informs the design of more efficient algorithms tailored to the presence of ties.
Beyond fixed-confidence BAI, related problems such as fixed-budget best-arm identification also face challenges with multiple optimal arms. Studies like (gabillon2012best, ) and (chen2014combinatorial, ) investigate how limited sampling budgets affect the ability to identify top arms accurately, with recent insights showing that tie-aware sampling can improve error rates. Furthermore, problems involving combinatorial bandits (chen2014combinatorial, ), contextual bandits (simchi2021best, ), and reinforcement learning where multiple optimal policies exist also benefit from algorithms capable of recognizing and exploiting ties among arms or actions.
Besides efficiency considerations, the presence of multiple optimal arms has practical implications for fairness and diversity. When several arms represent different treatments, products, or policies with equivalent expected performance, selecting any one of them ensures equitable treatment and avoids unintended bias in recommendations or decisions (xu2021fairness, ). Algorithms that identify ties quickly can facilitate fairer deployment strategies in real-world applications.
In summary, while much of the classical best-arm identification literature focuses on unique optima, recent research increasingly recognizes the importance of handling multiple optimal arms. Our work contributes to this growing area by refining the Track-and-Stop algorithm through a new stopping condition that is both tie-aware and theoretically grounded. We prove that this modified version achieves instance-optimal sample complexity in the fixed-confidence setting, even when the set of optimal arms is not a singleton. Our approach closes a fundamental theoretical gap in the literature and offers new insights into designing efficient, fair, and practically relevant algorithms for modern stochastic bandit problems.
@@ I-C Contributions
This work makes the following key contributions:
  • Modified Track-and-Stop with Optimal Stopping: We propose a refined version of the Track-and-Stop algorithm featuring a novel stopping rule that achieves instance-optimal sample complexity for best-arm identification under the fixed-confidence setting, even when multiple optimal arms exist. This resolves a long-standing open problem in the literature.
  • New Information-Theoretic Lower Bound: We introduce a novel lower bound on sample complexity that explicitly accounts for multiple optimal arms, characterizing the minimal number of samples required to identify any optimal arm with high confidence—avoiding unnecessary distinctions between tied arms.
  • Refined Theoretical Analysis: We develop an analytical framework that combines generalized likelihood ratio statistics with convex optimization techniques, enabling a tight alignment between the proposed stopping criterion and the information-theoretic lower bounds.
  • Adaptivity to Multiple Optimal Arms: Our analysis covers settings with multiple optimal arms, showing that the modified Track-and-Stop algorithm automatically adapts to such structures and preserves optimal performance guarantees.
@@ I-D Notations
For two distributions $P$ and $Q$ defined over a common support $\mathcal{X}$, the Kullback–Leibler (KL) divergence from $P$ to $Q$ is given by
[…]
@@
For any positive real number $x\in\mathbb{R}_{+}$, we use $\log(x)$ to denote the logarithm base 2. Furthermore, define the binary KL divergence function between two Bernoulli parameters $x,y\in[0,1]$ as
$\mathrm{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y},$
@@
For any functions $f:\mathbb{R}\to\mathbb{R}$, we define
$\dot{f}(x)=\frac{df(x)}{dx},\qquad\ddot{f}(x)=\frac{d^{2}f(x)}{dx^{2}}.$
Finally, let
$\Sigma_{K}=\left\{w\in\mathbb{R}_{+}^{K}:\sum_{i=1}^{K}w_{i}=1\right\}$
denote the $K$-dimensional probability simplex.
@@ II One-parameter exponential family
$p_{\mu}(x)=\exp\left(\theta(\mu)x-A(\theta(\mu))\right),$ (1)
where:
@@
$f(\lambda_{1},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
@@ III Problem Setup
Challenge with Multiple Optima.
@@ IV General Lower Bound
First, we recall the following lemma.
@@
(Kaufmann2014OnTC, , Lemma 1) Let $\nu$ and $\nu^{\prime}$ be two bandit models with $K$ arms such that for all $a$, the distributions $\nu_{a}$ and $\nu^{\prime}_{a}$ are mutually absolutely continuous. For any almost-surely finite stopping time $\sigma$ with respect to $(\mathcal{F}_{t})$,
@@
Denote
$\mathcal{S}=\big{\{}\mu:\mu_{1}=\mu_{2}=\cdots=\mu_{M}>\mu_{M+1}\geq\mu_{M+2}\geq\cdots\geq\mu_{K}\big{\}}.$ (10)
@@
For any tuple $(\alpha_{1},\alpha_{2},\cdots,\alpha_{M})\in[0,1]^{M}$ such that $0\leq\sum_{i=1}^{M}\alpha_{i}\leq 1$, define
$I_{\alpha_{1},\alpha_{2},\cdots,\alpha_{M}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a})$ $=\sum_{i=1}^{M}\alpha_{i}d\bigg{(}\mu_{i},\sum_{i=1}^{M}\alpha_{i}\mu_{i}+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}\mu_{a}\bigg{)}$
$\qquad+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}d\bigg{(}\mu_{a},\sum_{i=1}^{M}\alpha_{i}\mu_{i}+\bigg{(}1-\sum_{i=1}^{M}\alpha_{i}\bigg{)}\mu_{a}\bigg{)}.$ (11)
Let $\delta\in(0,1)$. For any $\delta$-PAC strategy and any bandit model $\mu\in\mathcal{S}$,
$\mathbb{E}_{\mu}[\tau_{\delta}]\geq T^{*}(\mu)\text{kl}(1-\delta\,\|\,\delta),$ (12)
@@
$T^{*}(\mu)^{-1}=\sup_{w\in\Sigma_{K}}\min_{a\notin\{1,2,\cdots,M\}}\bigg{(}\sum_{i=1}^{M}w_{i}+w_{a}\bigg{)}$
×$\qquad\times I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\frac{w_{2}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a}).$ (13)
@@
As $\text{kl}(1-\delta\,\|\,\delta)\sim\log(1/\delta)$ when $\delta$ goes to zero. Theorem 4 provides an asymptotic lower bound
$\liminf_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau_{\delta}]}{\log(1/\delta)}\geq T^{*}(\mu).$ (14)
A non-asymptotic version can be obtained by using the inequality $\text{kl}(1-\delta\,\|\,\delta)\geq\log(1/(2.4\delta))$ that holds for any $\delta\in(0,1)$ (Garivier2016OptimalBA, ).
@@
Denote by
$\text{Alt}(\mu)=\bigcup_{a\in\mathcal{A}\setminus\{1,2,\cdots,M\}}\big{\{}\lambda\in\mathcal{S}:(\lambda_{a}>\lambda_{1})\wedge(\lambda_{a}>\lambda_{2})\wedge\cdots\wedge(\lambda_{a}>\lambda_{M})\}.$ (15)
For any $\lambda\in\text{Alt}(\mu)$, there exists $a\notin\{1,2,\cdots,M\}$ such that $\lambda_{a}>\lambda_{i}$ for all $i\in[M]$. Hence, $\{1,2,\cdots,M\}$ is not the set of optimal arms under $\lambda$. Thus, introducing the event =$\mathcal{E}=\{\hat{a}_{\tau}\in\{1,2,\cdots,M\}\}\in\mathcal{F}_{\tau}$, any $\delta$-PAC algorithm satisfies $\mathbb{P}_{\mu}(\mathcal{E})>1-\delta$ and $\mathbb{P}_{\lambda}(\mathcal{E})\leq\delta$.
@@
$\forall\lambda\in\text{Alt}(\mu):\sum_{a=1}^{K}d(\mu_{a},\lambda_{a})\mathbb{E}_{\mu}[N_{a}(\tau_{\delta})]\geq\text{kl}(1-\delta,\delta).$ (16)
@@
$\text{kl}(\delta,1-\delta)\leq\mathbb{E}_{\mu}[\tau_{\delta}]\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg{)},$ (17)
@@
$\mathbb{E}_{\mu}[\tau_{\delta}]\geq\bigg{[}\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg{(}\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg{)}\bigg{]}^{-1}\text{kl}(1-\delta,\delta).$ (18)
@@
$\qquad=\min_{a\notin\{1,2,\cdots,M\}}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a}).$ (19)
Next, we minimize $f(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$ under constraints $\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}$. This is a convex optimization problem (see Lemma 2) that can solved analytically.
@@
$L(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a},\beta_{1},\beta_{2},\cdots,\beta_{M})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}),$ (20)
@@
where $\mu=A^{\prime}(\theta)$, the KL divergence between $P_{\mu}$ and $P_{\lambda}$ is given by:
@@
$\qquad=\sum_{i=1}^{M}w_{i}\big{[}A(\theta(\lambda_{i}))-A(\theta(\mu_{i}))-(\theta(\lambda_{i})-\theta(\mu_{i}))\mu_{i}\big{]}$
$\qquad\qquad+w_{a}\big{[}A(\theta(\lambda_{a}))-A(\theta(\mu_{a}))-(\theta(\lambda_{a})-\theta(\mu_{a}))\mu_{a}\big{]}-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}).$ (21)
@@
00 0 $=\frac{\partial L}{\partial\lambda_{i}}=w_{i}(\lambda_{i}-\mu_{i})\frac{d\theta(\lambda_{i})}{d\lambda_{i}}+\beta_{i},\qquad\forall i\in[M],$ (22)
00 0 $=\frac{\partial L}{\partial\lambda_{a}}=w_{a}(\lambda_{a}-\mu_{a})\frac{d\theta(\lambda_{a})}{d\lambda_{a}}-\sum_{i=1}^{M}\beta_{i},$ (23)
@@
Now, we consider cases:

### v2

@@ Abstract
We study best-arm identification in stochastic multi-armed bandits under the fixed-confidence setting, focusing on instances with multiple optimal arms. Unlike prior work that addresses the unknown-number-of-optimal-arms case, we consider the setting where the number of optimal arms is known in advance. We derive a new information-theoretic lower bound on the expected sample complexity that leverages this structural knowledge and is strictly tighter than previous bounds. Building on the Track-and-Stop algorithm, we propose a modified, tie-aware stopping rule and prove that it achieves asymptotic instance-optimality, matching the new lower bound. Our results provide the first formal guarantee of optimality for Track-and-Stop in multi-optimal settings with known cardinality, offering both theoretical insights and practical guidance for efficiently identifying any optimal arm.
@@ I Introduction
The multi-armed bandit (MAB) problem provides a fundamental framework for sequential decision-making under uncertainty, where a learner interacts with a set of arms corresponding to unknown reward distributions. The objective is to efficiently balance exploration and exploitation in order to make statistically reliable decisions.
In the best-arm identification (BAI) problem under the fixed-confidence setting, the goal is to identify an arm with the highest (or nearly highest) expected reward with a prescribed confidence level, using as few samples as possible. This formulation arises in various applications such as clinical trials, A/B testing, hyperparameter optimization, and recommendation systems.
Most existing works on BAI assume the existence of a unique best arm. However, in many real-world scenarios, multiple optimal arms may exist, each achieving the same maximal expected reward. Degenne and Koolen [1] studied this case and derived a fundamental lower bound together with the sticky Track-and-Stop algorithm, which attains asymptotic sample complexity matching this bound when the number of optimal arms is unknown. Despite these advances, the optimal strategy and the fundamental performance limits when the number of optimal arms is known a priori remain unexplored.
@@ I-A Motivation
The best-arm identification problem is central to sequential decision-making, where the objective is to identify an optimal arm using as few samples as possible while ensuring a fixed level of confidence. This problem arises in a wide range of domains, including clinical trials, A/B testing, and hyperparameter tuning. While the case of a unique best arm has been extensively studied, many practical applications naturally involve multiple arms achieving the same maximal expected reward. This motivates the need to understand the optimal sample complexity in this more general multi-optimal scenario.
Among existing algorithms for fixed-confidence BAI, the Track-and-Stop algorithm stands out for its principled foundation based on likelihood ratio statistics and its strong empirical performance. Although it is widely regarded as near-optimal in the single-optimum case, a complete theoretical justification of its optimality when multiple optimal arms exist has remained elusive.
Degenne and Koolen [1] addressed the setting where the number of optimal arms is unknown, establishing a lower bound and proposing the sticky Track-and-Stop algorithm, which asymptotically achieves this bound. In contrast, this paper focuses on the complementary setting where the number of optimal arms is known in advance. Under this assumption, we derive a new information-theoretic lower bound that characterizes the minimal achievable sample complexity, which is strictly tighter than the bound in [1].
Furthermore, we propose a modified version of the Track-and-Stop algorithm and prove that it achieves instance-optimal sample complexity in the fixed-confidence regime. These results provide the first formal characterization of the optimal sample complexity when multiple optimal arms exist and their number is known, thereby completing the theoretical picture and offering new insights into the design of exploration strategies in stochastic bandit problems.
@@ I-B Related Work
Best-arm identification (BAI) in stochastic multi-armed bandits has been a central focus in the literature for over a decade, particularly under fixed-confidence and fixed-budget settings. In the fixed-confidence framework, the goal is to identify an arm with the highest expected reward with probability at least $1-\delta$, while minimizing the expected number of samples [2, 3]. The fixed-budget setting, by contrast, fixes the total sampling budget and seeks to minimize the probability of misidentification [4].
@@
Early work introduced elimination-based strategies and stopping rules that guarantee PAC (Probably Approximately Correct) bounds [2], later complemented by information-theoretic approaches establishing matching lower and upper bounds under the assumption of a unique best arm [3]. Such analyses leverage Kullback–Leibler (KL) divergences to characterize instance difficulty, and algorithms like Track-and-Stop [3] and lil’UCB [5] exploit these insights for near-optimal performance. Other PAC-based algorithms, including LUCB [6], adapt confidence bounds to efficiently balance exploration, but rely on uniqueness of the best arm and may over-sample tied arms.
In practice, multiple arms often share the maximal expected reward, e.g., in clinical trials, recommendation systems, or A/B testing. Standard algorithms may waste samples distinguishing statistically equivalent arms, motivating approaches that exploit ties. Prior work has addressed related problems, such as top-k arm identification and best-subset selection [4, 7, 8], which aim to identify all optimal arms but often incur unnecessary sample complexity when the goal is to find any optimal arm. Degenne and Koolen [1] explicitly considered multiple optimal arms, deriving a lower bound and proposing the sticky Track-and-Stop algorithm, which achieves asymptotically optimal sample complexity when the number of optimal arms is unknown.
From a theoretical standpoint, classical lower bounds assume a unique best arm [3]; handling multiple optima requires refined bounds that characterize minimal sampling effort across tied arms. Beyond fixed-confidence BAI, related challenges arise in fixed-budget BAI [4, 7], combinatorial bandits [7], and contextual bandits [9], where tie-aware strategies improve both efficiency and fairness [10].
In summary, while most BAI studies focus on unique optima, recent work highlights the importance of efficiently handling multiple optimal arms. Our paper contributes by deriving a new lower bound for the case where the number of optimal arms is known and demonstrating that a modified Track-and-Stop algorithm achieves instance-optimal sample complexity, closing a key theoretical gap between the known- and unknown-cardinality settings.
@@ I-C Contributions
This paper makes the following contributions:
  • Tighter Fundamental Limit: We derive a new information-theoretic lower bound on the sample complexity of fixed-confidence BAI when the number of optimal arms is known, strictly improving over previous bounds for the unknown-cardinality setting [1].
  • Tie-Aware Track-and-Stop Algorithm: We propose a modification of the Track-and-Stop algorithm with a tie-aware stopping rule that leverages the known number of optimal arms to efficiently allocate samples among tied arms.
  • Instance-Optimality Guarantee: We prove that the modified algorithm achieves asymptotic instance-optimal sample complexity, matching the new lower bound and providing the first formal optimality guarantee for multi-optimal settings with known arm cardinality.
  • Insights for Efficient Exploration: Our results illustrate how structural knowledge of optimal arm multiplicity can be exploited to design exploration strategies that are both theoretically grounded and practically efficient.
@@ I-D Notations
For two probability distributions $P$ and $Q$ defined over a common support $\mathcal{X}$, the Kullback–Leibler (KL) divergence from $P$ to $Q$ is given by
$D_{\mathrm{KL}}(P\,\|\,Q)=\int_{\mathcal{X}}\log\!\left(\frac{dP}{dQ}(x)\right)dP(x),$ (1)
@@
For any positive real number $x\in\mathbb{R}_{+}$, we use $\log(x)$ to denote the logarithm base $2$. Furthermore, the binary KL divergence between two Bernoulli parameters $x,y\in[0,1]$ is defined as
$\mathrm{kl}(x,y)=x\log\frac{x}{y}+(1-x)\log\frac{1-x}{1-y},$ (2)
@@
For any differentiable function $f:\mathbb{R}\to\mathbb{R}$, we denote its first and second derivatives by
$\dot{f}(x)=\frac{df(x)}{dx},\qquad\ddot{f}(x)=\frac{d^{2}f(x)}{dx^{2}}.$ (3)
We denote the $K$-dimensional probability simplex by
$\Sigma_{K}=\left\{w\in\mathbb{R}_{+}^{K}:\sum_{i=1}^{K}w_{i}=1\right\}.$ (4)
Finally, for any $M\in\mathbb{N}^{+}$, we define the index set $M]=\{1,2,\ldots,M\}$.
@@ II One-parameter exponential family
$p_{\mu}(x)=h(x)\exp\left(\theta(\mu)x-A(\theta(\mu))\right),$ (5)
where $h(x)$ is a non-negative function, and
@@
$f(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$
@@ III Problem Setup
Challenge with Multiple Optima
@@ IV General Lower Bound
Let $N_{a}(t)=\sum_{s=1}^{t}\mathbf{1}_{\{A_{s}=a\}}$ be the number of draws of arm $a$ between the instants $1$ and $t$, and $N_{a}=N_{a}(\tau)$ be the total number of draws of arm $a$. First, we recall the following lemma.
@@
[11, Lemma 1] Let $\nu$ and $\nu^{\prime}$ be two bandit models with $K$ arms such that for all $a$, the distributions $\nu_{a}$ and $\nu^{\prime}_{a}$ are mutually absolutely continuous. For any almost-surely finite stopping time $\sigma$ with respect to $(\mathcal{F}_{t})$,
@@
Now, denote by
$\mathcal{S}=\big\{\mu=(\mu_{1},\mu_{2},\cdots,\mu_{K}):\mbox{exactly $M$ optimal arms}\big\}.$ (14)
Definition 4.
Let $a\in\{1,2,\cdots,K\}\setminus\{1,2,\cdots,M\}$. For any tuple $(\alpha_{1},\alpha_{2},\cdots,\alpha_{M})\in[0,1]^{M}$ such that $0\leq\sum_{i=1}^{M}\alpha_{i}\leq 1$, define
$I_{\alpha_{1},\alpha_{2},\cdots,\alpha_{M}}(\nu_{1},\nu_{2},\cdots,\nu_{M},\nu_{a})$
$\quad=\sum_{i=1}^{M}\alpha_{i}d\bigg(\nu_{i},\sum_{i=1}^{M}\alpha_{i}\nu_{i}+\bigg(1-\sum_{i=1}^{M}\alpha_{i}\bigg)\nu_{a}\bigg)$
$\qquad+\bigg(1-\sum_{i=1}^{M}\alpha_{i}\bigg)d\bigg(\nu_{a},\sum_{i=1}^{M}\alpha_{i}\nu_{i}+\bigg(1-\sum_{i=1}^{M}\alpha_{i}\bigg)\nu_{a}\bigg),$
$\qquad\forall(\nu_{1},\nu_{2},\cdots,\nu_{M},\nu_{a})\in\mathbb{R}^{M+1}.$ (15)
@@
Consider a bandit $\mu=(\mu_{1},\mu_{2},\cdots,\mu_{K})\in\mathcal{S}$ such that the set of optimal arms is $\mu_{1},\mu_{2},\cdots,\mu_{M}$. Let
$\text{Alt}(\mu)$ $=\bigcup_{a\in\mathcal{A}\setminus\{1,2,\cdots,M\}}\big\{\lambda\in\mathcal{S}:(\lambda_{a}>\lambda_{1})\wedge(\lambda_{a}>\lambda_{2})$
$\qquad\wedge\cdots\wedge(\lambda_{a}>\lambda_{M})\}.$ (16)
Let $\delta\in(0,1)$. Then, for any $\delta$-PAC strategy, it holds that
$\mathbb{E}_{\mu}[\tau]\geq T^{*}(\mu)\text{kl}(1-\delta,\delta),$ (17)
@@
$T^{*}(\mu)^{-1}=\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg(\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg)$ (18)
$=\sup_{w\in\Sigma_{K}}\min_{a\notin[M]}\bigg(\sum_{i=1}^{M}w_{i}+w_{a}\bigg)$
$\times I_{\frac{w_{1}}{\sum_{i=1}^{M}w_{i}+w_{a}},\frac{w_{2}}{\sum_{i=1}^{M}w_{i}+w_{a}},\cdots,\frac{w_{M}}{\sum_{i=1}^{M}w_{i}+w_{a}}}(\mu_{1},\mu_{2},\cdots,\mu_{M},\mu_{a}).$ (19)
@@
Some remarks are in order.
  • The case $M=K$ is trivial when $M$ is known in advance, as any arm can be selected. Hence, in this work we assume $K>M$. Moreover, Theorem 5 recovers the sample complexity result in [3, Lemma 3] when $M=1$.
  • As $\text{kl}(1-\delta,\delta)\sim\log(1/\delta)$ when $\delta$ goes to zero. Theorem 5 provides an asymptotic lower bound
    $\liminf_{\delta\to 0}\frac{\mathbb{E}_{\mu}[\tau]}{\log(1/\delta)}\geq T^{*}(\mu).$ (20)
    A non-asymptotic version can be obtained by using the inequality $\text{kl}(1-\delta,\delta)\geq\log(1/(2.4\delta))$ that holds for any $\delta\in(0,1)$ [3].
@@
For any $\lambda\in\text{Alt}(\mu)$, there exists $a\notin[M]$ such that $\lambda_{a}>\lambda_{i}$ for all $i\in[M]$. Hence, $[M]$ is not the set of optimal arms under $\lambda$. Thus, introducing the event $\mathcal{E}=\{\hat{a}_{\tau}\in[M]\}\in\mathcal{F}_{\tau}$, any $\delta$-PAC algorithm satisfies $\mathbb{P}_{\mu}(\mathcal{E})>1-\delta$ and $\mathbb{P}_{\lambda}(\mathcal{E})\leq\delta$.
@@
$\forall\lambda\in\text{Alt}(\mu):\sum_{a=1}^{K}d(\mu_{a},\lambda_{a})\mathbb{E}_{\mu}[N_{a}(\tau)]\geq\text{kl}(1-\delta,\delta).$ (21)
@@
$\text{kl}(1-\delta,\delta)\leq\mathbb{E}_{\mu}[\tau]\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg(\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg),$ (22)
@@
$\mathbb{E}_{\mu}[\tau]\geq\bigg[\sup_{w\in\Sigma_{K}}\inf_{\lambda\in\text{Alt}(\mu)}\bigg(\sum_{a=1}^{K}w_{a}d(\mu_{a},\lambda_{a})\bigg)\bigg]^{-1}\text{kl}(1-\delta,\delta).$ (23)
@@
$\qquad=\min_{a\notin[M]}\inf_{\lambda:\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}}\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})$
$\qquad\qquad+w_{a}d(\mu_{a},\lambda_{a}).$ (24)
Next, we minimize $f(\lambda_{1},\lambda_{2},\cdots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})$ under constraints $\lambda_{a}\geq\lambda_{1},\lambda_{a}\geq\lambda_{2},\cdots,\lambda_{a}\geq\lambda_{M}$. This is a convex optimization problem (see Lemma 2) that can solved analytically. First, we assume that $w_{i}>0,\forall i\in[M]$.
@@
$\quad=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}),$ (25)
@@
where $\mu=\dot{A}(\theta)$, the KL divergence between $P_{\mu}$ and $P_{\lambda}$ is given by:
@@
$\quad=\sum_{i=1}^{M}w_{i}\big[A(\theta(\lambda_{i}))-A(\theta(\mu_{i}))-(\theta(\lambda_{i})-\theta(\mu_{i}))\mu_{i}\big]$
$\quad\qquad+w_{a}\big[A(\theta(\lambda_{a}))-A(\theta(\mu_{a}))-(\theta(\lambda_{a})-\theta(\mu_{a}))\mu_{a}\big]$
$\qquad\qquad-\sum_{i=1}^{M}\beta_{i}(\lambda_{a}-\lambda_{i}).$ (26)
@@
$0=\frac{\partial L}{\partial\lambda_{i}}=w_{i}(\lambda_{i}-\mu_{i})\frac{d\theta(\lambda_{i})}{d\lambda_{i}}+\beta_{i},\qquad\forall i\in[M],$ (27)
$0=\frac{\partial L}{\partial\lambda_{a}}=w_{a}(\lambda_{a}-\mu_{a})\frac{d\theta(\lambda_{a})}{d\lambda_{a}}-\sum_{i=1}^{M}\beta_{i},$ (28)
@@
If $w_{a}=0$, then from (28) we obtain $\beta_{i}=0$ for all $i\in[M]$. Consequently, by (27), it follows that $\lambda_{i}=\mu_{i}$ for all $i\in[M]$. In this case, we have
$f(\lambda_{1},\lambda_{2},\ldots,\lambda_{M},\lambda_{a})=\sum_{i=1}^{M}w_{i}d(\mu_{i},\lambda_{i})+w_{a}d(\mu_{a},\lambda_{a})=0.$ (31)
Therefore, without loss of generality, we may assume that $w_{a}>0$.
We consider three cases: